OpenStudy (anonymous):

use rational exponents to simplify. sqrt(5)x^10

OpenStudy (saifoo.khan):

x^2

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt[5]{x^{10}}=x^{\frac{10}{5}}=x^2\] So \[\Large \sqrt[5]{x^{10}}=x^2\]

OpenStudy (anonymous):

cool! thats what i got! I like it when I can come up with the right answer even when I'm not sure i'm doing correctly!

OpenStudy (saifoo.khan):

=D